# Pump-efficient Josephson parametric amplifiers with high saturation power

### v1

@@ III.2.3 Bandwidth and IP3
Bandwidth is an important measure of amplifier performance due to the need to amplify signals over a range of frequencies. For the amplifiers in subsection III.2.1, we find a bandwidth of 279 MHz. The large bandwidth is a consequence of the high damping rate of about 3 GHz. We plot the gain of the nondegenerate amplifier over a range of frequencies in Fig. 17(a), demonstrating this bandwidth. The gain versus signal frequency plot for the nondegenerate amplifier described in subsection III.2.2 is shown in Fig. 17(b), with a bandwidth of 97 MHz.
@@
Another measure of amplifier performance is IP3, which indicates the range of signal power over which intermodulation products between two distinct signal frequencies remain small. Here, we compute the output power of third order intermodulation products resulting from the amplification of two similar but distinct signal frequencies within the amplifier bandwidth [23, p. 511-518]. We note that in order to reduce computation run time, we slightly modify the procedure in subsection II.3 to use the last half of the DTI solution rather than the last quarter, while ensuring the integration time remains at least 2000 periods of the signal frequencies. This allows us to choose an integration time which is twice $T_{\text{min}}$ rather than 4 times $T_{\text{min}}$. This reduced integration time does not affect our solution since for the two similar signal frequencies used in IP3 analysis, the integration time is very long and transient solutions die out far before half the integration time.
@@
We provide a diagram of the output power of our two signals as well as third order products in Fig. 18(a), which demonstrates the IP3 performance of the amplifier in subsection III.2.1. As anticipated, the output signal power grows linearly with input signal power, while the third order terms grow cubically until the saturation point. We find IIP3 of $-79.2$ dBm and OIP3 of $-62.3$ dBm. Similarly, in Fig. 18(b), we provide the IP3 performance of the amplifier in subsection III.2.2, where we find IIP3 of $-88.7$ dBm and OIP3 of $-75.7$ dBm.
@@
Figure 17: Amplifier bandwidth: gain of the amplifier versus signal frequency. a) Amplifier in subsection III.2.1 b) Amplifier in subsection III.2.2.
@@
Figure 18: IP3: output power of signal frequencies as well as their third order intermodulation products, plotted in dBm units. Here we have $\omega_{1}=101/201\omega_{\text{p}}$ in blue and $\omega_{2}=101/200\omega_{\text{p}}$ in orange. The third order term $2\omega_{1}-\omega_{2}$ is in green and $2\omega_{2}-\omega_{1}$ is in red. We include linear fits for $\omega_{2}$ and $2\omega_{1}-\omega_{2}$ which demonstrate the intercept points. These fits take into account the 3rd through 10th points of each data set. a) Amplifier in subsection III.2.1 b) Amplifier in subsection III.2.2.
III.3 Chained polynomial amplifiers
@@
Polynomial amplifiers always have a higher PAE at relatively low target gains (as shown in Fig. 7). Intuitively, without introducing extra higher order nonlinearity, we can obtain a higher PAE by connecting amplifiers with smaller gains as opposed to a single amplifier with higher gain. In this section, we first calculate the total PAE for a chained amplifier consisting of two amplifiers. Next, we use the result from Sec. III.1 to show that we can increase the total PAE significantly by tuning the gain of the two amplifiers. Finally, we extend our investigation to a chain of $n$ amplifiers that all have the same gain and PAE. These amplifiers are not identical, as $A_{\text{sat}}$ needs to be appropriately scaled to match the signal amplitude in the amplifier chain. By connecting approximately 30 amplifiers in a chain, we can achieve a PAE of 95% for a 20 dB degenerate amplifier.
@@
Figure 19: A schematic of an amplifier chain. Circulators and low-pass filters are added to route input, output, and pump waves.
@@
Figure 20: Total PAE of a chained 20 dB degenerate amplifier consists of two 7-th order polynomial amplifiers.
In subsection III.1, we optimized a range of polynomial amplifiers designed for various target gains. Specifically, for a target gain of $G_{\text{t}}=20$ dB, the highest PAE we obtained was $48.4\%$. To improve this efficiency, we connect a pair of smaller gain amplifiers to achieve a total gain of 20 dB with overall threshold ±$\pm 1$ dB. To characterize the amplifier chain, we used the PAE versus gain curve from Fig. 7. To ensure the overall gain variation does not break the ±$\pm 1$ dB threshold, we made sure that each amplifier in the chain does not break a threshold of ±$\pm 0.5$ dB. The performance of this amplifier chain is illustrated in Fig. 20. We observe that the total PAE of the chain peaks at 72% when the first amplifier is set to a gain of 13 dB and the second to a gain of 7 dB.
@@
Figure 21: Total PAE of a chained 20 dB degenerate amplifier consisting of $n$ same-performance amplifiers.
@@ IV Discussion
Due to the limited power added efficiency of modern JPAs and the need for low noise, efficient amplifiers in quantum computing applications, we have explored optimizing JPA designs by finely tuning the amplifier’s Hamiltonian. We find that with the freedom to define an arbitrary inductive block, we can design amplifiers with a PAE orders of magnitude higher than that of modern JPAs. Using polynomial amplifier designs, we find that we can reach a PAE of nearly 50% at 20 dB gain, which is the upper limit in the nondegenerate case.
Furthermore, restricting ourselves to inductive blocks consisting of circuits extending previous RF-SQUID designs, we still realize a significant improvement in PAE. We are able to design JPA circuits with high PAEs, large bandwidth, and limited intermodulation distortion. We also provide designs for amplifiers with reduced PAE which may prove to be more practically buildable. These amplifiers could prove to be useful in quantum computing applications which require the amplifier to be located in proximity to qubits at millikelvin temperatures, where minimizing heat dissipation is a priority. We anticipate many, or at least several, of these devices to be mounted in a single dilution refrigerator. Recent implementations of low-PAE, high-saturation power amplifiers are problematic [22, 28] due to the extremely high pump powers utilized in these experiments, which in turn generate excessive heat in the cryostat.
@@ Appendix B Harmonic Balance Method
We start with the degenerate polynomial amplifier, where $\omega_{\text{p}}=2\omega_{\text{s}}=2\omega_{0}$. In this case, we propose that the solution takes the following form:
$\phi(t)=a_{0}+a_{1}e^{i\omega_{0}t}+\cdots+a_{n}e^{in\omega_{0}t}+h.c..$ (29)
@@
where $\hat{D}$ is the differential operator, represented by the matrix $\hat{D}_{ab}=ia\omega_{0}\delta_{ab}$; $\mathbf{v}_{\text{in}}$ denotes the Fourier coefficients of the incoming wave; and $\mathcal{J}(\mathbf{v})$ is the nonlinear function in the frequency domain obtained via a Fourier transform of $J(\phi(t))$. The explicit form of $\mathcal{J}(\mathbf{v})$ is hard to find. Thus, we calculate $\mathcal{J}(\mathbf{v})$ using the Fourier transform,
@@
The situation becomes more complicated in the nondegenerate case, where $\omega_{\text{p}}=2\omega_{0}$ and $\omega_{\text{s}}=\omega_{0}\pm\delta\omega$. In this case, we need to add more harmonics in our proposed solutions. When the signal amplitude is much smaller than the pump amplitude, we only need to add the single signal photon process, i.e., the term with a frequency of $N\omega_{\text{p}}\pm\omega_{\text{s}}$. Then, we can use the same techniques to solve this nondegenerate amplifier. However, we need to be cautious that, in the Fourier transform (31), we apply proper zero-padding since we only keep the single signal photon process. Thus, the HB method is usually slower in the nondegenerate case. When the amplitude of the signal is comparable to that of the pump, the algorithm’s performance deteriorates as the higher order photon processes we neglect become significant. In principle, we could include additional components in the solution, but this will considerably slow down the HB algorithm.
Appendix C Numerical Solution and Optimization of JPA Circuits
@@
In solving the EOMs for a circuit of the form described in Fig. 1, we use the DTI method described in Section II.3 to compute the output signal from the amplifier. We use a total integration time that is commensurate with the period of the pump and signal frequencies. That is, we enforce that the integration time must be a multiple of $T_{\min}=\text{gcd}(\frac{2\pi}{\omega_{\text{p}}},\frac{2\pi}{\omega_{\text{s}}})$. We choose an integration time of $4n_{c}T_{\text{min}}$ since we wish to isolate the steady-state solution to the equations of motion, while at time near $t=0$ we see a transient solution to the differential equations. We choose $n_{c}$ such that the integration time represents at least 2000 periods of both the signal and pump. We take only the last quarter of the DTI solution, which we find is sufficient to remove transient effects from the output signal. In order to extract the power of the amplified signal frequency, we use the DFT with 50 samples of the DTI solution per signal period.

### v2

@@ III.3.3 Bandwidth and IP3
Bandwidth is an important measure of amplifier performance due to the need to amplify signals over a range of frequencies. For the amplifiers in Subsec. III.3.1, we find a bandwidth of 327 MHz. The large bandwidth is a consequence of the high damping rate of about $2.5~{}(2\pi)$ GHz. We plot the gain of the nondegenerate amplifier over a range of frequencies in Fig. 17(a), demonstrating this bandwidth. The gain versus signal frequency plot for the nondegenerate amplifier described in Subsec. III.3.2 is shown in Fig. 17(b), with a bandwidth of 375 MHz.
@@
Another measure of amplifier performance is IP3, which indicates the range of signal power over which intermodulation products between two distinct signal frequencies remain small. Here, we compute the output power of third order intermodulation products resulting from the amplification of two similar but distinct signal frequencies within the amplifier bandwidth, as is specified on pages 511-518 of Ref. [27]. We note that in order to reduce computation run time, we slightly modify the procedure in Subsec. II.3 to use the last half of the DTI solution rather than the last quarter, while ensuring the integration time remains at least 2000 periods of the signal frequencies. This allows us to choose an integration time which is twice $T_{\text{min}}$ rather than 4 times $T_{\text{min}}$. This reduced integration time does not affect our solution since for the two similar signal frequencies used in IP3 analysis, the integration time is very long and transient solutions die out far before half the integration time.
@@
We provide a diagram of the output power of our two signals as well as third order products in Fig. 18(a), which demonstrates the IP3 performance of the amplifier in Subsec. III.3.1. As anticipated, the output signal power grows linearly with input signal power, while the third order terms grow cubically until the saturation point. We find input third order intercept point (IIP3) of $-95.0$ dBm and output third order intercept point (OIP3) of $-78.9$ dBm. Similarly, in Fig. 18(b), we provide the IP3 performance of the amplifier in Subsec. III.3.2, where we find IIP3 of $-99.2$ dBm and OIP3 of $-82.2$ dBm. In each diagram, the black vertical line represents the input saturation power, demonstrating that the third order intermodulation products remain substantially smaller than the output signals up to this point. For further analysis of intermodulation products and a full power spectrum near the saturation point, see Appendix E.
@@
Figure 17: Amplifier bandwidth: gain of the nondegenerate amplifier versus signal frequency. (a) Amplifier in Subsec. III.3.1. (b) Amplifier in Subsec. III.3.2.
@@
Figure 18: IP3: output power of signal frequencies as well as their third order intermodulation products, plotted in dBm units. Input signal power refers to the total input power from both signals. Here we have $\omega_{1}=101/201\,\omega_{\text{p}}$ in blue and $\omega_{2}=101/200\,\omega_{\text{p}}$ in orange. The third order term $2\omega_{1}-\omega_{2}$ is in green and $2\omega_{2}-\omega_{1}$ is in red. We include fits for $\omega_{2}$ (dashed orange, which is the lower of the linear components), and $2\omega_{2}-\omega_{1}$ (dashed red, panel (a), which is the higher of the cubic components) and $2\omega_{1}-\omega_{2}$ (dashed green, panel (b), which is the higher of the cubic components) which demonstrate the intercept points. These fits take into account the 3rd through 10th points of each data set. The black lines represent the input saturation power as determined by our requirement that the input signal be amplified by 20 dB $\pm$ 1 dB. (a) Amplifier in Subsec. III.3.1. (b) Amplifier in Subsec. III.3.2.
III.4 Polynomial amplifier chains
@@
Polynomial amplifiers always have a higher PAE at relatively low target gains (as shown in Fig. 7). Intuitively, without introducing extra higher order nonlinearity, we can obtain a higher PAE by connecting amplifiers with smaller gains as opposed to a single amplifier with higher gain. In this section, we first calculate the total PAE for an amplifier chain consisting of two amplifiers. Next, we use the result from Subsec. III.1 to show that we can increase the total PAE significantly by tuning the gain of the two amplifiers. Finally, we extend our investigation to a chain of $n$ amplifiers that all have the same gain and PAE. These amplifiers are not identical, as $A_{\text{sat}}$ needs to be appropriately scaled to match the signal amplitude in the amplifier chain. By connecting 15 amplifiers in a chain, we can achieve a PAE of 98% for a 20 dB degenerate amplifier.
@@
Figure 19: A schematic of an amplifier chain. On the left hand side, we connect two amplifiers with gains of $G_{1}$ and $G_{2}$ and PAEs of $\eta_{1}$ and $\eta_{2}$, respectively. This amplifier chain is equivalent to a single amplifier with total gain $G_{\text{tot}}$ and achieves a total PAE of $\eta_{\text{tot}}$ shown on the right hand side. Perfect circulators and low-pass filters are added to route input, output, and pump waves.
@@
Figure 20: Total PAE of a 20 dB degenerate amplifier chain composed of two 6-th order polynomial amplifiers.
In Subsec. III.1, we optimized a range of $6$-th order polynomial amplifiers designed for various target gains. Specifically, for a target gain of $G_{\text{t}}=20$ dB, the highest PAE we obtained was $63.3\%$. To improve this efficiency, we connect a pair of smaller gain amplifiers to achieve a total gain of 20 dB, where we tolerate variations of up to ±$\pm 1$ dB from this target. To characterize the amplifier chain, we used the PAE versus gain curve from Fig. 7. To ensure the overall gain variation does not exceed the ±$\pm 1$ dB tolerance, we made sure that each amplifier in the chain does not vary from its target gain by more than ±$\pm 0.5$ dB. The performance of this amplifier chain is illustrated in Fig. 20. We observe that the total PAE of the chain peaks at 88.3% when the first amplifier is set to a gain of 12 dB and the second to a gain of 8 dB.
@@
Figure 21: Total PAE of a 20 dB degenerate amplifier chain composed of $n$ same-performance amplifiers.
@@ IV Discussion
Due to the limited power added efficiency of modern JPAs and the need for low noise, efficient amplifiers in quantum computing applications, we have explored optimizing JPA designs by finely tuning the amplifier’s inductive block. We find that with the freedom to define an arbitrary inductive block, we can design amplifiers with a PAE orders of magnitude higher than that of modern JPAs. In particular, modern JPAs currently in use are tuned so as to optimize third order nonlinearity. These intrinsic third order terms dynamically generate higher order nonlinearities, which shift signal and idler modes off resonance [25]. For example, the $g_{3}$ term in Eq. (22) generates a response at zero frequency and twice the signal frequency which is quadratic in signal amplitude. This in turn generates a cubic response at the signal frequency, which looks like the $g_{4}$ term of Eq. (22). Here, we counter this effect by tuning intrinsic higher order nonlinearities in order to cancel the dynamically generated nonlinearities. Using polynomial amplifier designs, we find that we can reach a PAE of around 63% at 20 dB gain, which is the upper limit in the nondegenerate case.
Restricting ourselves to inductive blocks composed of circuit elements, we identify a rule for designing JPAs. In particular, we find that the inductive block should have a monotonic current-phase relation to avoid issues with how phase divides among elements in the block. Further, considering inductive blocks consisting of circuits extending previous RF-SQUID designs with monotonic current-phase relations, we still realize a significant improvement in PAE. We are able to design JPA circuits with high PAE, large bandwidth, and limited intermodulation distortion. In particular, we provide designs for amplifiers based on chains of RF-SQUIDs with high PAE, and provide designs with chains of extended RF-SQUIDs with even higher PAE at the expense of manufacturing complexity. In Appendix F, we discuss the robustness of these designs to small manufacturing imperfections.
These amplifiers could prove to be useful in quantum computing applications which require the amplifier to be located in proximity to qubits at millikelvin temperatures, where minimizing heat dissipation is a priority. Compared to previous JPA designs, the amplifiers in this paper are able to provide similar output signal power while requiring a substantially weaker pump, thereby reducing heat load. These high saturation power and high PAE amplifiers may be particularly advantageous in scenarios where simultaneous multiplexed qubit readout is necessary [^3] Our amplifiers continue to demonstrate good PAE when amplifying multiple distinct signals. Appendix E addresses the power spectrum for two signals near the saturation point., as the power requirement scales with the number of qubits. Additionally, these amplifiers demonstrate low distortion at low signal power, which provides a significant benefit in using high saturation power amplifiers as compared to low saturation power amplifiers [34, 24]. We anticipate many, or at least several, of these devices to be mounted in a single dilution refrigerator. Recent implementations of low PAE, high saturation power amplifiers are problematic [26, 32] due to the extremely high pump powers utilized in these experiments, which in turn generate excessive heat in the cryostat.
@@ Appendix B Harmonic balance method
We start with the degenerate polynomial amplifier, where $\omega_{\text{p}}=2\omega_{\text{s}}$. In this case, we propose that the solution takes the following form:
$\phi(t)=a_{0}+a_{1}e^{i\omega_{\text{s}}t}+\cdots+a_{n}e^{in\omega_{\text{s}}t}+h.c..$ (34)
@@
where $\hat{D}$ is the differential operator, represented by the matrix $\hat{D}_{ab}=ia\omega_{\text{s}}\delta_{ab}$; $\mathbf{v}_{\text{in}}$ denotes the Fourier coefficients of the incoming wave; and $\mathcal{J}(\mathbf{v})$ is the nonlinear function in the frequency domain obtained via a Fourier transform of $J(\phi(t))$. The explicit form of $\mathcal{J}(\mathbf{v})$ is hard to find. Thus, we calculate $\mathcal{J}(\mathbf{v})$ using the Fourier transform,
@@
The situation becomes more complicated in the nondegenerate case, where $\omega_{\text{p}}=\omega_{\text{s}}+\omega_{\text{i}}$ and $\delta\omega=(\omega_{\text{s}}-\omega_{\text{i}})/2\neq 0$. In this case, we need to add more harmonics in our proposed solutions. When the signal amplitude is much smaller than the pump amplitude, we only need to add the single signal photon process, i.e., the term with a frequency of $N\omega_{\text{p}}\pm\omega_{\text{s}}$. Then, we can use the same techniques to solve this nondegenerate amplifier. However, we need to be cautious that, in the Fourier transform (36), we apply proper zero-padding since we only keep the single signal photon process. Thus, the HB method is usually slower in the nondegenerate case. When the amplitude of the signal is comparable to that of the pump, the algorithm’s performance deteriorates as the higher order photon processes we neglect become significant. In principle, we could include additional components in the solution, but this will considerably slow down the HB algorithm.
Appendix C Avoiding local minima in energy-phase relations and enforcing monotonicity in current-phase relations
In this work, we found that certain amplifier designs produce unstable amplification. As a result, we consider two different limitations on amplifier design which we must enforce when optimizing our amplifiers. First, for all amplifiers of the form in Fig. 1, it is crucial that $\phi=0$ is the unique minimum of the energy function $E[\phi]$. This condition ensures that the amplifier does not jump between different local minima as the signal amplitude increases, which could result in unreliable amplification. Since current is the derivative of energy with respect to phase, we impose this condition by requiring that the current is strictly positive for $\phi>0$ and strictly negative for $\phi<0$. Second, amplifiers with inductive blocks with identical repeating elements must have monotonic current-phase relations, as discussed in Subsec. III.2. This requirement is stricter than the first requirement.
To enforce these requirements while optimizing our amplifiers, we introduce a penalty term in the cost function to discourage designs that violate these constraints. Since it is impractical to track all current values during the optimization process, we instead generate a discrete set of points from the current-phase relation within the range $[\Phi_{-},\Phi_{+}]$ with an appropriate density. This phase range must be sufficiently large to account for fluctuations in $\phi(t)$ throughout the process. For polynomial amplifiers, to enforce that the current is strictly negative for $\phi<0$ and strictly positive for $\phi>0$, we define the following penalty function:
$\text{penalty}=\lambda\big{(}\max(0,I^{\max}_{-})+\max(0,-I^{\min}_{+})\big{)},$ (38)
where $\lambda$ is the regularization factor, and
$I^{\max}_{-}=\max\big{(}\{I(\phi)|\Phi_{-}\leq\phi<0\}\big{)},$ $I^{\min}_{+}=\min\big{(}\{I(\phi)|0<\phi\leq\Phi_{+}\}\big{)}.$ (39)
For inductive blocks composed of identical repeating elements such as our extended RF-SQUIDs, we check that the current increases between any two successive points with increasing phase, guaranteeing monotonicity. In this case, for each extended RF-SQUID, we consider a grid of 100 points for phase values between $-6\pi$ and $6\pi$. The penalty which enforces this is
$\text{penalty}=\lambda\big{(}\tanh(-\Delta\text{I}_{\text{min}}-0.1)+1\big{)},$ (40)
where $\Delta\text{I}_{\text{min}}$ is the minimum difference between sequential current values in the grid.
In the case of an inductive block with repeating RF-SQUIDs, we impose this requirement analytically, which is more computationally efficient. Considering the EOM in Eq. (18), the current-phase relation is monotonic when the oscillations of the $\sin$ term are smaller than the slope of the linear term. That is, we enforce that $i_{\text{c}}L/\phi_{0}<1$. In this case, the penalty is
$\text{penalty}=\lambda\big{(}\tanh(~{}20~{}(i_{\text{c}}L/\phi_{0}-1.15))+1\big{)}.$ (41)
For JPA circuits, we also find that it can be useful to use an alternative cost function for finer optimization from a high-PAE starting point. In these cases, we use the cost function
$f\left(\omega_{0},K,\cdots\right)=\sum_{m=1}^{N}\exp{\left(-1500/\left(G\left(A_{\text{s}}=\frac{mR}{N},\omega_{0},K,\cdots\right)-G_{\text{t}}\right)^{2}\right)}+\text{penalty}\end{aligned}$ (42)
instead of the cost function specified in Eq. (23). Here, if $G=G_{t}$ for any value of $m$, then the cost for that point is 0.
Appendix D Numerical solution and optimization of JPA circuits
@@
In solving the EOMs for a circuit of the form described in Fig. 1, we use the DTI method described in Section II.3 to compute the output signal from the amplifier. We use a total integration time that is commensurate with the period of the pump and signal frequencies. That is, we enforce that the integration time must be a multiple of $T_{\min}=\text{lcm}(\frac{2\pi}{\omega_{\text{p}}},\frac{2\pi}{\omega_{\text{s}}})$. We choose an integration time of $4n_{c}T_{\text{min}}$ since we wish to isolate the steady-state solution to the equations of motion, while at time near $t=0$ we see a transient solution to the differential equations. We choose $n_{c}$ such that the integration time represents at least 2000 periods of both the signal and pump. We take only the last quarter of the DTI solution, which we find is sufficient to remove transient effects from the output signal. To see that the transient solution quickly dissipates, we provide examples of the time-domain output signal computed via DTI for amplifiers with inductive blocks composed of a chain of 10 RF-SQUIDs. These examples, shown in Fig. 22, demonstrates that the transient solution gives way to the steady-state in the long-time regime for amplifiers with inductive blocks with monotonic current-phase relations. For those with nonmonotonic current-phase relations, the integrated solution does not become periodic. Since we only select the last quarter of our DTI solution, and only propose amplifiers with inductive blocks monotonic current-phase relations, we avoid including transient effects in analyzing the output of these amplifiers. In order to extract the power of the amplified signal frequency, we use the DFT with 50 samples of the DTI solution per signal period.
Figure 22: Example time-domain output signals computed by DTI for amplifiers with 10 RF-SQUIDs. In each panel, we plot the signal for short-time, demonstrating the transient behavior, and for long-time, demonstrating the steady-state behavior. (a) Amplifier with inductive block whose current-phase relation is very monotonic, i.e. its slope is never near zero. In this case, we see the initial transient behavior, and then the steady state behavior, with agreement among simulations with and without stray capacitances, and at intermediate phases. (b) Amplifier with inductive block whose current-phase relation is nearly nonmonotonic, i.e. it is monotonic but its slope nears zero. Here, we see transient behavior, but the steady state settles with agreement among simulations. (c) Amplifier with inductive block whose current-phase relation is nonmonotonic. In this case, the amplifier never settles, behavior is not periodic even in the long-time regime, and the simulations do not agree.
@@
Appendix E Power Spectrum of Nondegenerate Amplifier Near Gain Compression Point
Here we investigate the power spectrum of the nondegenerate RF-SQUID amplifier when used to amplify two signals near the saturation point. This serves to identify the behavior of other intermodulation products in addition to the third order terms which are relevant in IP3 analysis. We provide the output power spectrum of the band with focus on the frequencies near the signal and pump tones, as well as $\omega_{\text{p}}+\omega_{\text{s}}$, in Fig. 23. That is, we can investigate intermodulation products around 6 GHz, 12 GHz, and 18 GHz.
The highest peak in the spectrum is the pump tone at 12 GHz, while the two signals and their idlers on either side of 6 GHz follow. The next largest components, the third order products plotted in Fig. 18 and the second order terms at $\omega_{\text{p}}+\omega_{\text{s}}$, remain a factor of at least 21 below the two signal tones. We therefore believe intermodulation products remain manageable up to saturation power, and we note that the saturation point is reached not because of pump power being consumed by these intermodulation products, but because of fundamental limits to the amplifier’s ability to amplify at the signal frequency.
Figure 23: Fourier analysis of the output of the nondegenerate RF-SQUID amplifier near the saturation point. Here we have the two signal frequencies $\omega_{1}=101/201\,\omega_{\text{p}}$ and $\omega_{2}=101/200\,\omega_{\text{p}}$, which are given the same input power (-105.9 dBm). The top portion of the figure shows the spectrum of the full band, from 0 GHz to 25 GHz. We provide further detail of the power spectrum around 6 GHz, 12 GHz, and 18 GHz, which shows the vicinity of the signal frequencies and the pump frequency, as well as second order terms which mix pump and signal. The two peaks just above 6 GHz represent the two signals which are approximately 1 to 2 orders of magnitude higher than the next highest products (except their idlers, which are approximately equal), and are lower than only the pump tone at 12 GHz.
Appendix F Tolerance of amplifier performance to variations in circuit parameters
The JPA circuit designs provided in this paper and their predicted PAE are derived by considering ideal circuits, where repeating blocks are able to be manufactured identically and with no variation in parameters. We consider here how the performance of the degenerate RF-SQUID amplifier specified in Table 2 is affected by variations in the fluxes through each of the 10 loops in our circuit. To do this, we simulate the circuit with intermediate nodes and stray capacitances as in Sec. III.2. In particular, we simulated the RF-SQUID amplifier circuit with 10 equations of motion, where each equation corresponds to a single RF-SQUID, and the flux through each SQUID’s loop is varied. Since this is a computationally difficult problem, we computed amplifier performance for a single set of fluxes which varied from the optimized flux by a random amount selected via a normal distribution. We increased the standard deviation of this distribution to test the effect of increasing variations on amplifier performance.
Figure 24: Gain versus signal amplitude for different amounts of variation in the external flux through each RF-SQUID loop. In the low signal amplitude regime, increasing the variation increases gain, leaving the desired ±$\pm 1~{}\text{dB}$ region. For higher signal amplitudes, the performance tends to be less sensitive to variations in flux.
We find that when adding variations to the flux through the loop, we must adjust the pump power to maintain our 20 dB target gain in the low signal regime. In Fig. 24, we show gain versus signal amplitude for varying standard deviation of the fluxes without adjusting pump power. The gain at low signal amplitude quickly leaves the 20 dB $\pm$ 1dB region, as indicated by the black horizontal lines.
Figure 25: Gain versus signal amplitude for select amounts of variation in external flux. We reduce signal power gradually to re-enter the ±$\pm 1~{}\text{dB}$ region. (a) Variations with standard deviation of $0.003\phi_{0}$. (b) Variations with standard deviation of $0.005\phi_{0}$. (c) Variations with standard deviation of $0.01\phi_{0}$.
We then consider decreasing the pump power for different amounts of flux variation to return to our 20 dB $\pm$ 1dB regime and recompute PAE. We find that for a standard deviation of $0.003\phi_{0}$, we are able to retain a PAE of at least 34.3% (Fig. 25(a)), as compared to 37.9% without the added flux variations. That is, with this level of variation in flux, we lose only a marginal amount of efficiency. Increasing the standard deviation to $0.005\phi_{0}$, we find a PAE of at least 11.3% (Fig. 25(b)). For a standard deviation of $0.01\phi_{0}$, the PAE is at least 8% (Fig. 25(c)). Therefore, for this JPA circuit design, it is ideal to keep variations in threaded flux below $0.003\phi_{0}$ to $0.005\phi_{0}$.
Another possible manufacturing defect is a uniform offset of a device parameter from the ideal value described in this paper. In practical use, the ideal operating point of an amplifier is determined by adjusting both pump power and frequency so as to maximize PAE. In order to demonstrate how this can mitigate uniform manufacturing defects, we consider the case of our degenerate RF-SQUID circuit specified in Table 2 with a decrease of 1%, 3% and 5% of the linear inductances from their ideal values. We adjust the pump power and signal/pump frequency together in order to achieve the highest possible PAE under these defects, as compared to 37.9% with no defects. With 1% offset, we find that the PAE can be restored to at least 33.0%, with 3% offset, we can achieve at least 20.8% PAE, and with 5% offset, we can achieve at least 9.2% PAE. We provide the gain versus input signal power in these three cases in Fig. 26. We therefore find that it is possible to recover amplifier performance in practical amplifiers which are subject to the limitations of manufacturing precision.
Figure 26: Gain versus input signal power for 1%, 3%, and 5% decreases in the linear inductances in the degenerate RF-SQUID amplifier. The pump power and pump/signal frequencies have been adjusted to recover PAE. As compared to the original design, in the 1% case the pump and signal frequencies are increased by 1.4% and the pump power is -72.2 dBm. For the 3% offset, the frequencies are increased by 3.7% and the pump power is -71.9 dBm. For the 5% offset, the frequencies are increased by 5% and the pump power is -71.5 dBm.